OpenStudy (anonymous):

Solve the combined inequality and select the answer that best describes the graph of the solution. 4x + 6 < 14 or 8x - 6 > 26

OpenStudy (anonymous):

My answer is: 8x - (4x + 6) = 14 4x =14 +6 4x =20 x =20/4 x =5 <(^.^)>

OpenStudy (anonymous):

Thanks :3

OpenStudy (anonymous):

Your Welcome!